OpenStudy (anonymous):

Which of the following shows a pair of equations that are equivalent? A. 8 + 4(x + y) = 12 4(2 + x + y) = 12 B. 8 + 4(x + y) = 12 8 + 4x + 4y = 48 C. 8 + 4(x + y) = 12 8 + 4x + y = 12 D. 8 + 4(x + y) = 12 12(x + y) = 12

OpenStudy (anonymous):

can someone help pleeeze

OpenStudy (anonymous):

A. 8 + 4(x + y) = 12 8 + 4x + 4y = 12 4x + 4y = 12 - 8 4x + 4y = 4 4(2 + x + y) = 12 8 + 4x + 4y = 12 4x + 4y = 12 - 8 4x + 4y = 4 ANSWER : A